myininaya (myininaya):

satellite, i can't get back to the problem i had 8a+2b=-1 and 10a+2b=3

myininaya (myininaya):

no need to solve lol

OpenStudy (smurfy14):

a=2 b=-17/2

OpenStudy (anonymous):

yeah me neither. i got locked out

myininaya (myininaya):

i was just telling smurfy so he wouldn't waste his time lol

OpenStudy (anonymous):

i had the second equation but not the first. but i forget what i had. i forget the problem. and you are probably right and i am wrong

OpenStudy (smurfy14):

oh :/ lol

OpenStudy (anonymous):

the first part set f(x)=4 when you plug in 2 that is all i remember

myininaya (myininaya):

but gj smurfy

myininaya (myininaya):

can you open my attachments?

OpenStudy (anonymous):

any way i am sure you got rid of it on purpose so i could not do it. that's ok i don't care. nope. not at all

OpenStudy (smurfy14):

lol thanks

myininaya (myininaya):

lol you such a meanie

myininaya (myininaya):

always blaming me

OpenStudy (anonymous):

you are the bike hog, not me

myininaya (myininaya):

im sorry amistre likes me more

OpenStudy (anonymous):

yes you are more likable.

myininaya (myininaya):

don't be jealous

OpenStudy (smurfy14):

myininaya did u ever paint ur nails? lol

OpenStudy (anonymous):

ok now i got let you try to catch up. my new goal is half of amistre's medals.

myininaya (myininaya):

lol i did but i couldn't put the pi symbol on there it would cost me 7.00 dollars and i dont think pi is worth it

OpenStudy (anonymous):

(8a+2b=-1)*-1------>-8a-2b=1 10a+2b=3 -8a-2b=1 --------- 2a=4 a=2 8a+2b=-1 8(2)+2b=-1 16+2b=-1 2b=-1-16 2b=-17 b=-17/2

myininaya (myininaya):

lol no one has to solve this problem

OpenStudy (smurfy14):

lol

OpenStudy (anonymous):

ok but where did the 8a+2b=-1 come from?

myininaya (myininaya):

left limit of 2=4 right limit of 2=a2^3+2b+3=8a+2b+3 for actual limit to exist 4=8a+2b+3

OpenStudy (anonymous):

it is the given.

OpenStudy (anonymous):

hmmmmmmmm

myininaya (myininaya):

lol pat

OpenStudy (anonymous):

then shouldnt it be 8a+ 2 b = 1 like i said?!!

OpenStudy (anonymous):

and not, for example, -1?

myininaya (myininaya):

omg omg no! are you for real did i screw up i screwed up

OpenStudy (anonymous):

i will try really hard not to gloat

OpenStudy (anonymous):

I think you made a mistake in the first equation myininaya. I found that \(a=\frac{1}{3}\) and \(b=-\frac{1}{6}\).

myininaya (myininaya):

i wrote down 8a+2b+3=2 instead of 8a+2b+3=4

OpenStudy (anonymous):

i get a = 1, but hey, maybe i am wrong....

OpenStudy (anonymous):

It should be \(4a+2b+3=4\).

OpenStudy (anonymous):

we need the original problem back

OpenStudy (anonymous):

just follow what i said.

myininaya (myininaya):

dang it i wrote ax^3+bx+3

myininaya (myininaya):

pat you don't know what we are talking about

OpenStudy (anonymous):

But strangely you got the 2nd equation right :D

OpenStudy (anonymous):

4a + 2b + 3 = 4 4a + 2b = 1 10a + 2b = 3 6a= 2 a= 1/3 yes?

myininaya (myininaya):

i did it so quickly and retartly i suck

OpenStudy (anonymous):

i am sorry.

OpenStudy (anonymous):

Yes a=1/3 and b=-1/6.

OpenStudy (anonymous):

this has been better entertainment that i deserve. goodnight all

myininaya (myininaya):

i will give you a medal pat for caring

myininaya (myininaya):

he likes to see me fail

OpenStudy (anonymous):

who?!

myininaya (myininaya):

satellite

OpenStudy (anonymous):

haha really? why is that?

myininaya (myininaya):

ask him something about a bike i dont remember

OpenStudy (anonymous):

maybe you always beat him in answering questions :P

myininaya (myininaya):

lol maybe

myininaya (myininaya):

i think i need to go to bed you still have a good brain anwar at nighttime

OpenStudy (anonymous):

It's 7:23 AM where I am right now :D

myininaya (myininaya):

did galaxy get the right answers?

OpenStudy (anonymous):

I wrote the answer there. I wish he/she already got it.

myininaya (myininaya):

he could probably figure out if i made any mistakes anyways if he knows alittle bit about what he is doing

OpenStudy (anonymous):

Yeah!

myininaya (myininaya):

goodnight anwar

OpenStudy (anonymous):

Good night!